# Cost-aware Photonic Graph State Generation: A Graphical Framework

### v1

@@ Abstract
Photonic graph states are essential resources for quantum computation and communication. Deterministic emitter-based generation of graph states overcomes the scalability issues of probabilistic approaches, but their experimental realization is constrained by technological demands, often expressed by the number of two-qubit gates and the depth and/or width of the quantum circuits used to model the generation process. We introduce a cost-aware framework for the generation of photonic graph states of arbitrary size and shape, built on a complete set of necessary and sufficient conditions and a universal set of elementary graph operations that govern the evolution of the state toward the target. Within this framework, we develop Graph Builder, a deterministic generation algorithm that achieves substantial reductions—up to an order of magnitude—in two-qubit gate usage for both random and structured graphs, compared with alternative approaches. The algorithm uses the minimum number of emitters possible for a fixed emission sequence, while also supporting the use of extra emitters for controlled trade-offs between emitter count and other cost metrics. Moreover, by systematically identifying the degrees of freedom at each stage of the generation process, this framework fully characterizes the optimization landscape, enabling analytic, heuristic, or exhaustive strategies for further cost reductions. Our approach provides a general and versatile tool for designing and optimizing emitter-based photonic graph state generation protocols, essential for scalable and resource-efficient photonic quantum information processing.
@@ I Introduction
Building on these tools, we present Graph Builder, an algorithm for generating arbitrary photonic graph states using deterministic emitter-based platforms. The algorithm achieves substantial reductions in the use of two-qubit gates for preparing random graphs, achieving average (maximum) reductions of up to 80% (90%) and 52% (70%) compared to the two alternative methods presented in Refs. [55] and [50] (both based on the algorithm introduced by Ref. [48]), respectively. We also show consistent improvements in the use of two-qubit gates in preparation of structured graphs, such as Raussendorf–Harrington–Goyal lattices [56, 57], tree graphs [58], and parity encoded 6-ring states [6]. We remark that the reported improvements arise from the inherent cost-awareness built into the algorithm’s design, enabled by the new framework, rather than from dedicated heuristic or exhaustive optimization efforts. In fact, further cost reductions are possible by leveraging the relevant degrees of freedom available in the algorithm. As a key feature, the introduced algorithm enables a modular optimization framework in which explicit cost functions can be assigned to each decision point whenever there is freedom in the choice of operations. This facilitates the implementation of analytic, exhaustive, or heuristic optimization strategies that can be applied on top of the base generation algorithm to obtain further improvements. The usefulness of such secondary optimizations has been demonstrated in previous works [49, 50, 51].
@@
Lastly, while our algorithm has the advantage of operating with the minimum number of emitters required to generate a target photonic graph state under a fixed emission order, this constraint can be relaxed. Prior work [45] has shown that introducing additional emitters during generation can reduce other costs, such as gate count and circuit depth, which may be more relevant for certain experimental platforms. To support such trade-offs, we provide a straightforward generalization of our algorithm that accommodates extra emitters when the minimum-emitter constraint is lifted.
@@ II.3 Local Complementation
    Employing either of the updating methods shown in equations 4 and 5 leads to the same final adjacency matrix as the initial matrix is always symmetric.
@@ III The generative set
$G^{\prime}(0)\xrightarrow{\mathcal{T}}\ldots G^{\prime}(n)\ldots\xrightarrow{\mathcal{T}}G^{\prime}(N)=G.$ (8)
@@ III.1 Eligibility Conditions
$S_{\mathcal{P}}=-{\text{Tr}\hskip 2.84526pt}\left(\rho_{\mathcal{P}}\log\rho_{\mathcal{P}}\right),$ (9)
@@
$S_{\mathcal{P}}(G^{\prime})\geq S_{\mathcal{P}}(G),$ (10)
@@
$\mathrm{rank}[B(n)]=m^{\prime}\leq\min\{n,N-n\}.$ (11)
@@
$m^{\prime}\leq\mathrm{rank}[B^{\prime}(n)]\leq\min\{n,m\}.$ (12)
@@
  2. (II)
    Let $\mathcal{N}(i)=\{j\;|\;B^{\prime}_{ij}=1\}$ denote the set of indices corresponding to the emitters that are neighbors (connected) to photon $i$ in the physical graph. Then, for every $0\leq i<n$, the following must hold for the corresponding row $R_{i}$ in $B$:
    $R_{i}=\sum_{j\in\mathcal{N}(i)}R_{e_{j}}.$ (15)
@@
Conditions (I) and (II) only concern the bipartite entanglement across the inside-outside partition. We need to also ensure that the “inside” subsystem’s internal edge pattern at any step allows $G^{\prime}(n)$ to be a member of the GenSet($n$). As non-local interactions are forbidden between emitted photons, the changes that can be made in future steps in this internal entanglement structure are limited. Consequently, we need to add the following to the set of necessary conditions for $G^{\prime}(n)\in$ GenSet($n$) at step $n$:
  1. (III)
    For any graph $H$ and a subset of its nodes $\mathcal{V}$, let $H[\mathcal{V}]$ be the induced subgraph of $H$ on $\mathcal{V}$. Then for $\mathcal{P}=\{i\mid 0\leq i<n\}$, the subgraph $G^{\prime}[\mathcal{P}]$ must be equal to $\tilde{G}[\mathcal{P}]$ where $\tilde{G}$ is some member of the Partial Local Clifford equivalence class PLC($G$, $\mathcal{P}$) as defined below.
@@
$G[\mathcal{V^{\prime}}]\in PLC(G,\mathcal{V^{\prime}}),\quad\forall\;\mathcal{V}^{\prime}\subseteq V,$ (17)
@@
  • ($\text{III}^{\prime}$)
    For any intermediate physical graphs $G^{\prime}(n)$, we must have $G^{\prime}[\mathcal{P}]=G[\mathcal{P}]$ where $G$ is the target graph and $\mathcal{P}=\{i\mid 0\leq i<n\}$.
@@ IV.3 Emission Modes
  • Star (S) Mode: This emission adds a photonic node to the system, which is—upon emission—only connected to all neighbors of its emitter and not the emitter itself. This happens if we apply LC on the emitter node before and after a leaf emission.
@@ V Graph Builder Algorithm
We can now look back at our inductive algorithm; building upon the base case and utilizing the discussed eligibility conditions, and in particular eqs. 19 and 20, we provide the following requirements in accordance with the inductive step:
@@ VI.1 Random Graphs
The number of two-qubit gates utilized to generate random connected graphs for a range of graph sizes is provided in FIG. 14(a). Each data point represent the average value over a sample of 1000 random graphs of the same size with an edge density—the ratio of the number of present edges to the maximum possible number of edges in a graph—of 10%. Figure 14(b) shows the average and maximum of achieved reductions of our method relative to the alternative algorithms for each size. Average reductions of up to 80% and 52%, and maximum reductions of up to 90% and 70%, are observed relative to the time-reversed-1 and time-reversed-2 methods, respectively. It is worth noting that, on average, the number of two-qubit gates also depends on the edge density in random graphs. This behavior and performance as a function of edge density are analyzed in appendix C. The scaling of runtime with both size and edge density is also discussed in appendix D.
@@ VI.2.1 Encoded 6-Rings
Six-ring graphs have been proposed as resource states for fusion-based quantum computing [6]. Here, we consider the parity-encoded variant of these states, which offers higher loss thresholds for both computation and encoded-fusion processes [66, 67, 13]. The cnot preparation cost for an $(n,m)$-encoded six-ring graph, made up of $n\times m\times 6$ physical qubits (see FIG. 15) as defined in [68], depends only on the first encoding parameter $n$ and is given by $6n+4$ when using the our proposed algorithm, with only three emitters used in the generation process. In contrast, the alternative methods, time-reversed-1 and time-reversed-2, require 8n48n-4 and 18n1418n-14 cnots to prepare the same state, corresponding to overheads of 33% and 200%, respectively.
@@ VI.2.2 Branching Tree Cluster States
The branching tree cluster states are characterized by a branching vector $\vec{b}=\left(b_{1},\dots,b_{d}\right)$, where $b_{i}$ denotes the number of branches originating from each node at the $i$-th depth level. See FIG. 15 for an example. These states can be used as error correction codes when attached to qubits in other graph states, providing up to 50% loss tolerance [58], e.g., in a 2D cluster state for universal measurement-based quantum computing (MBQC). The code can also be used to reduce the failure probability of encoded-fusion operations, for instance in the generation of 3D fault-tolerant MBQC lattices [28]. For a homogeneous, depth $d$, branching vector where all $b_{i}$ are equal to a fixed branching parameter $b$, our algorithm generates the tree graph using $b^{d-1}-1$ cnot gates and $d$ emitters. As seen in Table 1, a cost comparison with alternative methods shows possible improvements of up to 50% and 20% compared with time-reversed-1 and time-reversed-2, for different depth and branching parameters.
@@
Tree (3,3,3) 3 40 39 8 10 20.0 16 50.0
Tree (4,4,4) 3 85 84 15 17 11.8 30 50.0
Tree (3,3,3,3) 4 121 120 26 33 21.2 56 53.6
Tree (4,4,4,4) 4 341 340 63 69 8.7 130 51.5
Tree (5,5,5,5) 4 781 780 124 144 13.9 252 50.8
Tree (3,3,3,3,3) 5 364 363 80 96 16.7 166 51.8
Tree (3,3,3,3,3,3) 6 1093 1092 242 291 16.8 492 50.8
RHG (1,1,1) 4 18 24 14 14 0.0 30 53.3
RHG (2,1,1) 4 31 44 28 29 3.4 68 58.8
RHG (3,1,1) 4 44 64 42 44 4.5 100 58.0
RHG (2,2,1) 7 53 80 56 61 8.2 130 56.9
RHG (3,2,1) 7 75 116 84 96 12.5 218 61.5
RHG (3,3,1) 10 106 168 126 150 16.0 302 58.3
RHG (2,2,2) 12 90 144 108 144 25.0 302 64.2
RHG (3,3,2) 17 179 300 240 369 35.0 804 70.1
RHG (3,3,3) 24 252 432 354 610 42.0 1300 72.8
@@ VI.2.3 RHG lattice
The RHG lattice [56, 57], designed for fault-tolerant MBQC [3, 4], consists of cubic unit cells with nodes located at the centers of the faces and edges of each cube (see Fig. 15). Here, we consider the $(L_{x},L_{y},L_{z})$-RHG lattice as defined in Ref. [68], where each parameter $L_{x,y,z}$ specifies the number of unit cells stacked along the corresponding axis. The cnot cost as the lattice size scales, comparing different algorithms can be found in Table 1. Improvements of up to 72% and 42% are achieved for the $(3,3,3)$-RHG state relative to the time-reversed-1 and time-reversed-2 algorithms, respectively.
@@ VI.2.4 Repeater Graph States
Figure 15: Examples for different classes of graphs. (a) Branching tree: {3, 3, 3}. (b) 16-node RGS. (c) 16-node generalized RGS. (d) RHG lattice unit cell. (e) 6-ring: (4, 2) parity encoded.
@@ VII.3.4 Local equivalence in intermediate steps
To make use of this freedom, let us look back at eq. 24:
$\mathcal{K}=\left\{\,e_{k}\;\middle|\;[R_{e_{k}}(n)]_{0}=1\right\},$
where $R]_{0}$ indicates the first element of the row vector $R$. The set $\mathcal{K}$ consists of emitters $\{e_{k}\}$ associated with the emitter rows $R_{e_{k}}$ which have non-zero first elements. In order for condition (III) to hold, the newly emitted photon at each step should get connected to photons in the neighborhood of emitters in $\mathcal{K}$. These connections are implemented by performing an e-to-inside operation between the new photon’s emitter and each $e_{k}\in\mathcal{K}$, and then emitting with an appropriate emission mode so that those edges transfer onto the newly emitted photon. Since each e-to-inside uses one cz gate, the two-qubit cost at step $n$ depends on $\lvert\mathcal{K}\rvert$. Now in an arbitrary step $n$, let us temporarily switch the target state from $G$ to $\tilde{G}$. Since at the beginning of the step we have $G^{\prime}(n)\in$ GenSet($G,n$), then the graph $\tilde{G}^{\prime}(n)\in$ GenSet($\tilde{G},n$) can be obtained from $G^{\prime}(n)$; this is because, by definition, $G$ can be converted to $\tilde{G}$ under local operation on all nodes and non-local gates on the nodes $\not\in$ $\mathcal{P}$. The same transformation can be applied on $G^{\prime}$ to convert it to $\tilde{G}^{\prime}$ where any Clifford operation applied on the nodes $\not\in$ $\mathcal{P}$ is replaced by a set of operation on the appropriate emitter nodes. Having in mind that emitter rows each correspond to one of the rows in $B(n)$, they are transformed accordingly when the target’s biadjacency matrix $B(n)$ is replaced by $\tilde{B}(n)$ of the PLC-equivalent graph $\tilde{G}$. The set $\mathcal{K}$ can now be presented as:
$\tilde{\mathcal{K}}=\left\{\,e_{k}\;\middle|\;[\tilde{R}_{e_{k}}(n)]_{0}=1\right\},$ (27)

### v2

@@ Abstract
Photonic graph states are essential resources for quantum computation and communication. Deterministic emitter-based generation of graph states overcomes the scalability issues of probabilistic approaches; nonetheless, their experimental realization is constrained by technological demands, often expressed by the number of two-qubit gates and the depth and/or width of the quantum circuits used to model the generation process. Here, we introduce a cost-aware framework for the generation of photonic graph states of arbitrary size and shape, built on a complete set of necessary and sufficient conditions and a universal set of elementary graph operations that govern the evolution of the state toward the target. Within this framework, we develop Graph Builder, a deterministic generation algorithm that achieves substantial reductions—up to an order of magnitude—in two-qubit gate usage for both random and structured graphs, compared with alternative approaches. Furthermore, we show that this framework enables the identification of elementary building blocks in specific cases, such as encoded 6-ring states. The algorithm uses the minimum number of emitters possible for a fixed emission sequence, while also supporting the use of extra emitters for controlled trade-offs between emitter count and other cost metrics. Moreover, by systematically identifying the degrees of freedom at each stage of the generation process, this framework fully characterizes the optimization landscape, enabling analytic, heuristic, or exhaustive strategies for further cost reductions. Our approach provides a general and versatile tool for designing and optimizing emitter-based photonic graph state generation protocols, essential for scalable and resource-efficient photonic quantum information processing.
@@ I Introduction
Building on these tools, we present Graph Builder, an algorithm for generating arbitrary photonic graph states using deterministic emitter-based platforms. The algorithm achieves substantial reductions in the use of two-qubit gates for preparing random graphs, achieving average (maximum) reductions of up to 61% (80%) and 52% (70%) compared to the two alternative methods presented in Refs. [40] and [58] (both based on the algorithm introduced by Ref. [38]), respectively. We also show consistent improvements in the use of two-qubit gates in preparation of structured graphs, such as Raussendorf–Harrington–Goyal lattices [49, 48], tree graphs [61], and parity-encoded 6-ring states [6], for which we also identify the explicit elementary generation circuit blocks. We remark that the reported improvements arise from the inherent cost-awareness built into the algorithm’s design, enabled by the new framework, rather than from dedicated heuristic or exhaustive optimization efforts. In fact, further cost reductions are possible by leveraging the relevant degrees of freedom available in the algorithm. As a key feature, the introduced algorithm enables a modular optimization framework in which explicit cost functions can be assigned to each decision point whenever there is freedom in the choice of operations. This facilitates the implementation of analytic, exhaustive, or heuristic optimization strategies that can be applied on top of the base generation algorithm to obtain further improvements. The usefulness of such secondary optimizations has been demonstrated in previous works [23, 58, 52].
@@
Lastly, while our algorithm has the advantage of operating with the minimum number of emitters required to generate a target photonic graph state under a fixed emission order, this constraint can be relaxed. Prior work [32] has shown that introducing additional emitters during generation can reduce other costs, such as gate count and circuit depth, which may be more relevant for certain experimental platforms. To support such trade-offs, we provide a straightforward generalization of our algorithm that accommodates extra emitters when the minimum-emitter constraint is lifted. We study the effect of using extra emitters on different cost metrics and observe reductions of up to 72% in circuit depth and 45% in two-qubit gate count for random graphs.
@@ II.3 Local Complementation
    Employing either of the updating methods shown in eqs. (4) and (5) leads to the same final adjacency matrix as the initial matrix is always symmetric.
@@ III The generative set
$G^{\prime}(0)\xrightarrow{\mathcal{T}}\ldots G^{\prime}(n)\ldots\xrightarrow{\mathcal{T}}G^{\prime}(N)=G$ (8)
@@ III.1 Eligibility Conditions
$S_{\mathcal{P}}=-{\text{Tr}\hskip 2.84526pt}\left(\rho_{\mathcal{P}}\log\rho_{\mathcal{P}}\right)$ (9)
@@
$S_{\mathcal{P}}(G^{\prime})\geq S_{\mathcal{P}}(G)$ (10)
@@
$\mathrm{rank}[B(n)]=m^{\prime}\leq\min\{n,N-n\}$ (11)
@@
$m^{\prime}\leq\mathrm{rank}[B^{\prime}(n)]\leq\min\{n,m\}$ (12)
@@
  2. Condition (II): Let $\mathcal{N}(i)=\{j\;|\;B^{\prime}_{ij}=1\}$ denote the set of indices corresponding to the emitters that are neighbors (connected) to photon $i$ in the physical graph. Then, for every $0\leq i<n$, the following must hold for the corresponding row $R_{i}$ in $B$:
    $R_{i}=\sum_{j\in\mathcal{N}(i)}R_{e_{j}}$ (16)
@@
Conditions (I) and (II) only concern the bipartite entanglement across the inside-outside partition. We need to also ensure that the “inside” subsystem’s internal edge pattern at any step allows $G^{\prime}(n)$ to be a member of the GenSet($n$). As non-local interactions are forbidden between emitted photons, the changes that can be made in future steps in this internal entanglement structure are limited. Consequently, we need to add the following to the set of necessary conditions for $G^{\prime}(n)\in$ GenSet($n$) at step $n$.
  1. Condition (III): For any graph $H$ and a subset of its nodes $\mathcal{V}$, let $H[\mathcal{V}]$ be the induced subgraph of $H$ on $\mathcal{V}$. Then for $\mathcal{P}=\{i\mid 0\leq i<n\}$, the subgraph $G^{\prime}[\mathcal{P}]$ must be equal to $\tilde{G}[\mathcal{P}]$ where $\tilde{G}$ is some member of the Partial Local Clifford equivalence class PLC($G$, $\mathcal{P}$) as defined below.
@@
$G[\mathcal{V^{\prime}}]\in PLC(G,\mathcal{V^{\prime}}),\quad\forall\;\mathcal{V}^{\prime}\subseteq V$ (18)
@@
  1. Condition (III): For any intermediate physical graphs $G^{\prime}(n)$, we must have $G^{\prime}[\mathcal{P}]=G[\mathcal{P}]$ where $G$ is the target graph and $\mathcal{P}=\{i\mid 0\leq i<n\}$.
@@ IV.3 Emission Modes
  • Star (S) Mode: This emission adds a photonic node to the system, which is—upon emission—only connected to all neighbors of its emitter and not the emitter itself. This happens if we apply LC on a neighbor of the emitter node, then apply LC on the emitter before a leaf emission, followed by LC on the emitter and LC on the same neighbor of the emitter.
@@ V Graph Builder Algorithm
We can now look back at our inductive algorithm; building upon the base case and utilizing the discussed eligibility conditions, and in particular eqs. (20) and (21), we provide the following requirements in accordance with the inductive step:
@@ VI.1 Random Graphs
The number of two-qubit gates utilized to generate random connected graphs for a range of graph sizes is provided in Fig. 14(a). Each data point represent the average value over a sample of 1000 random graphs of the same size with an edge density—the ratio of the number of present edges to the maximum possible number of edges in a graph—of 10%. Figure 14(b) shows the average and maximum of achieved reductions of our method relative to the alternative algorithms for each size. Average reductions of up to 61% and 52%, and maximum reductions of up to 80% and 70%, are observed relative to the time-reversed-1 and time-reversed-2 methods, respectively. It is worth noting that, on average, the number of two-qubit gates also depends on the edge density in random graphs. This behavior and performance as a function of edge density are analyzed in appendix C. The scaling of runtime with both size and edge density is also discussed in appendix D.
Note that the presented results are obtained using the minimum number of emitters possible to generate each graph. This constraint on the number of emitters used in the generation process can be relaxed to allow for potential reductions in other metrics of interest  [^2] For more details on this modification of the Graph Builder algorithm refer to section VII.3.5.. To showcase this, we have employed a greedy strategy to use a new emitter in each step of the generation whenever the introduction of the extra emitter results in a reduction in the number of cnot gates used in the current step that leads to the emission of the next photon. The outcome of such an analysis can be found in Fig. 15, where the generation circuit for a sample of random graphs were found once with the Graph Builder algorithm under the minimum emitter constraint and another time under the extra emitter strategy. A simultaneous reduction in the usage of the cnots and two-qubit depth of the generation circuit (e.g. the maximum number of two-qubit gates the emitters have to go through before being measured) is evident with an average (maximum) reduction of 3838% (7272%) in depth and 1818% (4545%) in CNOT gates for each case. The average of the number of emitters used for each case over the random sample was moved from 14 to 17, indicating that the mentioned advantages can be acquired at the cost of using approximately $3$ extra emitters in this case. For an explicit example of reducing the cnot cost by using additional emitters, see also section D of the supplementary material [1].
@@
Figure 15: Distribution of cnot gate count (left) and circuit depth (right) compared between using minimum emitters versus allowing for extra emitters over a sample of random graphs. A clear shift to the lower cost regime is visible for both metrics with a lower mean value and a tighter distribution when extra emitters are employed. The average number of emitters is increased from 14 to 17 between the two methods. The demonstrated results are for a sample of 1000 random 30-node graphs with a 20% edge density.
@@ VI.2 Special Cases
Figure 16: Examples for different classes of graphs. (a) 6-ring: (4, 2) parity encoded. (b) Branching tree: {3, 3, 3}. (c) RHG lattice unit cell. (d) 16-node RGS. (e) 16-node generalized RGS.
@@ VI.2.1 Encoded 6-Rings
Six-ring graphs have been proposed as resource states for fusion-based quantum computing [6]. Here, we consider the parity-encoded variant of these states, which offers higher loss thresholds for both computation and encoded-fusion processes [47, 36, 57]. We analyzed the cnot preparation cost for an $(n,m)$-encoded six-ring graph, as defined in Ref [35], made up of $n\times m\times 6$ physical qubits (see Fig. 16). As with many other parameterized classes of graphs, repeated patterns can emerge as the defining parameters are varied to scale the graph size. This allows one to identify circuit blocks in the generation circuit that repeat a number of times, depending on the parameter that controls the size of the graph. By inspecting this behavior as we change the values of $n$ and $m$, we find that the cnot cost depends only on the first encoding parameter $n$ and is given by $6n+4$ when using the our proposed algorithm, with only three emitters used in the generation process. In contrast, the alternative methods, time-reversed-1 and time-reversed-2, require 8n48n-4 and 9n79n-7 cnots to prepare the same state. This is confirmed by mathematical induction, as increasing $n$ to $n+1$ results in the addition of a circuit block with a certain number of cnot gates to the generation circuit of the new graph in each case, leading to the linear scaling with $n$.
The explicit form of such circuit blocks for each of three generation protocols, along with the numerical support for exact cnots relations can be found in section C of supplementary material [1].
Considering the asymptotic behavior for large $n$, which is of interest since the error-correction properties of the encoded graphs improve with increasing $n$ and $m$, we see that the $8n$ and $9n$ scaling of the alternative methods correspond to overheads of 33% and 50%, respectively when compared to the $6n$ case of the new algorithm.
@@ VI.2.2 Branching Tree Cluster States
The branching tree cluster states are characterized by a branching vector $\vec{b}=\left(b_{1},\dots,b_{d}\right)$, where $b_{i}$ denotes the number of branches originating from each node at the $i$-th depth level. See Fig. 16 for an example. These states can be used as error correction codes when attached to qubits in other graph states, providing up to 50% loss tolerance [61], e.g., in a 2D cluster state for universal measurement-based quantum computing (MBQC). The code can also be used to reduce the failure probability of encoded-fusion operations, for instance in the generation of 3D fault-tolerant MBQC lattices [39]. For a homogeneous, depth $d$, branching vector where all $b_{i}$ are equal to a fixed branching parameter $b$, our algorithm generates the tree graph using $b^{d-1}-1$ cnot gates and $d$ emitters. As seen in Table 1, a cost comparison with alternative methods shows possible improvements of up to 20% compared to time-reversed-2, for different depth and branching parameters.
@@
Tree (3,3,3) 3 40 39 8 10 20.0 8 0.0
Tree (4,4,4) 3 85 84 15 17 11.8 15 0.0
Tree (3,3,3,3) 4 121 120 26 33 21.2 28 7.1
Tree (4,4,4,4) 4 341 340 63 69 8.7 65 3.1
Tree (5,5,5,5) 4 781 780 124 144 13.9 126 1.6
Tree (3,3,3,3,3) 5 364 363 80 96 16.7 83 3.8
Tree (3,3,3,3,3,3) 6 1093 1092 242 291 16.8 246 1.6
RHG (1,1,1) 4 18 24 14 14 0.0 15 6.7
RHG (2,1,1) 4 31 44 28 29 3.4 34 17.6
RHG (3,1,1) 4 44 64 42 44 4.5 50 16.0
RHG (2,2,1) 7 53 80 56 61 8.2 65 13.8
RHG (3,2,1) 7 75 116 84 96 12.5 109 22.9
RHG (3,3,1) 10 106 168 126 150 16.0 151 16.6
RHG (2,2,2) 12 90 144 108 144 25.0 151 28.5
RHG (3,3,2) 17 179 300 240 369 35.0 402 40.3
RHG (3,3,3) 24 252 432 354 610 42.0 650 45.5
@@ VI.2.3 RHG lattice
The RHG lattice [49, 48], designed for fault-tolerant MBQC [50, 51], consists of cubic unit cells with nodes located at the centers of the faces and edges of each cube (see Fig. 16). Here, we consider the $(L_{x},L_{y},L_{z})$-RHG lattice as defined in Ref. [35], where each parameter $L_{x,y,z}$ specifies the number of unit cells stacked along the corresponding axis. The cnot cost as the lattice size scales, comparing different algorithms can be found in Table 1. Improvements of up to 45% and 42% are achieved for the $(3,3,3)$-RHG state relative to the time-reversed-1 and time-reversed-2 algorithms, respectively.
@@ VII.3.4 Local equivalence in intermediate steps
To make use of this freedom, let us look back at eq. 25, where $R]_{0}$ indicates the first element of the row vector $R$. The set $\mathcal{K}$ consists of emitters $\{e_{k}\}$ associated with the emitter rows $R_{e_{k}}$ which have non-zero first elements. In order for condition (III) to hold, the newly emitted photon at each step should get connected to photons in the neighborhood of emitters in $\mathcal{K}$. These connections are implemented by performing an e-to-inside operation between the new photon’s emitter and each $e_{k}\in\mathcal{K}$, and then emitting with an appropriate emission mode so that those edges transfer onto the newly emitted photon. Since each e-to-inside uses one cz gate, the two-qubit cost at step $n$ depends on $\lvert\mathcal{K}\rvert$. Now in an arbitrary step $n$, let us temporarily switch the target state from $G$ to $\tilde{G}$. Since at the beginning of the step we have $G^{\prime}(n)\in$ GenSet($G,n$), then the graph $\tilde{G}^{\prime}(n)\in$ GenSet($\tilde{G},n$) can be obtained from $G^{\prime}(n)$; this is because, by definition, $G$ can be converted to $\tilde{G}$ under local operation on all nodes and non-local gates on the nodes $\not\in$ $\mathcal{P}$. The same transformation can be applied on $G^{\prime}$ to convert it to $\tilde{G}^{\prime}$ where any Clifford operation applied on the nodes $\not\in$ $\mathcal{P}$ is replaced by a set of operation on the appropriate emitter nodes. Having in mind that emitter rows each correspond to one of the rows in $B(n)$, they are transformed accordingly when the target’s biadjacency matrix $B(n)$ is replaced by $\tilde{B}(n)$ of the PLC-equivalent graph $\tilde{G}$. The set $\mathcal{K}$ can now be presented as:
$\tilde{\mathcal{K}}=\left\{\,e_{k}\;\middle|\;[\tilde{R}_{e_{k}}(n)]_{0}=1\right\}$ (28)